OpenStudy (anonymous):

What is the perimeter ratio if the similarity ratio is 2:3 I would really appreciate help...

OpenStudy (raden):

easy. the ratio of perimeter still 2:3 too except, the question asking ratio of area. it is 2^2 : 3^2 = 4 : 9

OpenStudy (anonymous):

Great thank you. What is the similarity ratio if the perimeter ratio is 12:96?

OpenStudy (raden):

like i said before, similarity ratio = perimeter ratio so, still 12 : 96

OpenStudy (raden):

or simplied be 1 : 8

OpenStudy (anonymous):

So the. If the area ratio is 256:784, then the perimeter and similarity ratio are both 16:28?

OpenStudy (raden):

yes, that's right !

OpenStudy (anonymous):

Great. One more question: Two similar rhombus have areas of 125 and 245 . The perimeter of the smaller rhombus is 85. What is the perimeter of the larger rhombus? What is one side of the larger rhombus?

OpenStudy (anonymous):

Please help. That was the final question....

OpenStudy (raden):

retriceumed the perimeter of the larger is x. take the ratio : (perimeter smaller/perimeter larger)^2 = 125/245 (85/x)^2 = 125/245 (85/x)^2 = 25/49 take the square root on the right side, get 85/x = sqrt(25/49) 85/x = sqrt(25)/sqrt(49) 85/x = 5/7 cross multiplying, get 5x = 7 * 85 5x = 595 x = 595/5 x = 119 so, the perimeter of the larger that rhombus is 119

OpenStudy (raden):

opss, the 1st should *assumed

OpenStudy (anonymous):

What does that mean? And what is the side of the larger rhombus?

OpenStudy (anonymous):

So the large perimeter is 119?

OpenStudy (raden):

yes and then to get the side of the larger is just divided by 4 because all side of a rhombus is same. what is 119/4 ?

OpenStudy (anonymous):

29.75... Great thank you so much.merry late Christmas

OpenStudy (raden):

that's correct! you're welcome